The Ksp of PbSO4 is 1.3 x 10-8. What is the concentration of Pb2+ ions in a saturated solution of PbSO4?
Answer #1

given Ksp = 1.3 * 10^(-8)

PbSO4~~>Pb2+ + SO42-

PbSO4 concentration Pb2+ concentration SO42- concentration
initially. - say A 0 0
finally - A-x x x

Ksp = [Pb2+] [SO42-]

1.3 * 10^(-8) = x. x = x^(2)

x =(1.3 * 10^(-8))1/2

x = 1.14 * 10^(-4) M

thus concentration of Pb2+ is 1.14 *10^(-4) in saturated solution of PbSO4 which we calculated by using formula of Ksp.